OpenStudy (anonymous):

What is the value of y in the following system of equations? 2x+3y=6 2x+y=-2 thank you.

OpenStudy (anonymous):

2y =-8(by use the subtractoin) y=-4

OpenStudy (anonymous):

There isnt a negative four as an option. Only a positive four. ?

OpenStudy (diyadiya):

Subtract the two equations 2x+3y =6 2x+ y =-2 -------- 0 + 2y = 8 (3y-y=2y & 6-(-2) = 8 ) 2y=8 divide both sides by 2 y= ?

OpenStudy (anonymous):

okay sorry

OpenStudy (anonymous):

I divided 8 by 2 and got four. Thank you for your help and explaining.

OpenStudy (diyadiya):

@irishmolly You're Welcome :)